OpenStudy (anonymous):

A line has slope 1/4 and y-intercept 4. Which is the correct equation that describes this line in slope-intercept form?

OpenStudy (rulnick):

y = (1/4) x + 4

OpenStudy (anonymous):

y=x/4+4 or y=(1/4)x+4